OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x - 17 y = 3x - 5 y = 3x + 3

sam (.sam.):

\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

point slope form: y - 6 = 3(x - 1) Distribute it: y - 6 = 3x - 3 Solve for y and put into slope intercept form: + 6 y = 3x +3

sam (.sam.):

y-6=3(x-1) y=3x-3+6 y=3x+3